# FreeRide: Harvesting Bubbles in Pipeline Parallelism

### v1

@@ 8. Discussion
Security. Prior GPU sharing solutions tend to prioritize efficiency and assume a safe environment (yu_salus_2019, ; xiao_gandiva_2018, ; zhang_pilotfish_2022, ; lim_2021_zico, ; han_2022_reef, ). E.g., Orion assumes that the co-located GPU workloads are in the same trust domain (strati_orion_2024, ). FreeRide provides the same security and isolation guarantees as the lower-level system it is built upon. It uses MPS to limit GPU memory which provides separate GPU address spaces (mps_memory_protection, ) for pipeline training and side tasks, and Docker for environment isolation (docker_security, ; yasrab_2023_mitigating, ). Orthogonally, security for co-located GPU workloads is an active research area (pavlidakis_2024_gsafe, ; Naghibijouybari_2018_rendered, ; Naghibijouybari_2021_side, ; kim_2020_gpu, ; liu_2019_side, ; zhang_2024_bridge, ). We expect future work to co-design security with efficient GPU sharing.
Side task management. By implementing different strategies in its side task manager, FreeRide can incorporate more complicated management, e.g., co-locating multiple side tasks with various performance characteristics to improve the utilization of bubbles (veltair, ) or serving side tasks with fairness or performance guarantees (drf, ; chen_prophet_2017, ).
Scalability. FreeRide can be extended for better scalability. As FreeRide implements the communications among its components using RPCs, it can be easily extended to distributed settings with side tasks on multiple servers. FreeRide can also be extended for multi-GPU side tasks, e.g., distributed training and big data processing (mapreduce, ; ghive_socc, ), by launching workers with access to multiple GPUs.
Other ML accelerators. This work targets GPUs due to their widespread accessibility. FreeRide’s mitigation for bubbles fundamentally applies to other ML accelerators (tpu2017isca, ; mtiav1, ), provided that the platform has isolation and resource limit options for each process. We anticipate future work to incorporate the approach of FreeRide with other ML platforms.
@@ 9. Conclusion
We propose FreeRide, a system to harvest the bubbles in pipeline parallelism to serve generic GPU side tasks. It provides programming interfaces that abstract the life cycle of a side task as different states of a state machine and allows programmers to implement side tasks with little engineering effort. The side task manager and side task workers manage bubbles and side tasks and reduce the performance overhead of side tasks on pipeline training. Our evaluation shows that, on average, FreeRide achieves 7.8% cost savings for long-running, expensive, pipeline training with a negligible performance overhead of only about 1%.

### v2

@@ 8. Discussion
Security. Prior GPU sharing solutions tend to prioritize efficiency and assume a safe environment (yu_salus_2019, ; xiao_gandiva_2018, ; zhang_pilotfish_2022, ; lim_2021_zico, ; han_2022_reef, ). E.g., Orion assumes that the co-located GPU workloads are in the same trust domain (strati_orion_2024, ). FreeRide provides the same security and isolation guarantees as the lower-level system it is built on. It incorporates MPS to limit GPU memory which provides separate GPU address spaces (mps_memory_protection, ) for pipeline training and side tasks, and Docker for environment isolation (docker_security, ; yasrab_2023_mitigating, ). Orthogonally, security for co-located GPU workloads is an active research area (pavlidakis_2024_gsafe, ; Naghibijouybari_2018_rendered, ; Naghibijouybari_2021_side, ; kim_2020_gpu, ; liu_2019_side, ; zhang_2024_bridge, ). We expect future work to co-design security with efficient GPU sharing.
Fault tolerance. Since FreeRide supports generic side tasks, it is not possible for FreeRide to exhaustively implement fault tolerance mechanisms for all side tasks. Instead, FreeRide assumes that side tasks implement their own fault tolerance mechanisms to tolerate the failure of side tasks themselves and of pipeline training. In addition, since FreeRide deploys side tasks in Docker containers as processes that are independent of the pipeline training, failures of side tasks, such as illegal memory access, will not impact the main pipeline training workload.
Side task management. By implementing different strategies in its side task manager, FreeRide can incorporate more sophisticated management, e.g., co-locating multiple side tasks with various performance characteristics in the same worker to improve the utilization of bubbles (veltair, ) or serving side tasks with fairness or performance guarantees (drf, ; chen_prophet_2017, ).
Scalability. FreeRide can be extended for better scalability. As FreeRide implements communication among its components using RPCs, it can be easily extended to distributed settings with side tasks on multiple servers. During training, the side task manager of FreeRide receives bubbles from all GPUs from both remote servers and manages the side tasks that co-locate with each GPU. FreeRide can also be extended to support multi-GPU side tasks, e.g., distributed training and big data processing (mapreduce, ; ghive_socc, ), by launching workers with access to multiple GPUs.
Stability of pipeline training. FreeRide follows the same assumption as the previous pipeline parallel training works that pipeline training has a stable throughput and pattern, and that the training sequences have the same length after padding (liu_hanayo_2023, ; fan_dapple_2021, ; huang_gpipe_2019, ; qi_zero_2023, ).
Other ML accelerators. This work targets GPUs due to their widespread accessibility. FreeRide’s mitigation for bubbles fundamentally applies to other ML accelerators, such as Google’s TPU (tpu2017isca, ) and Meta’s MTIA (mtiav1, ), provided that the platform has isolation and resource limit options for each process. We anticipate future work to incorporate the approach of FreeRide with other ML platforms.
Energy consumption. There has been recent interest in building energy- and carbon-efficient systems for machine learning workloads (choi_envpipe_2023, ; ecoserve, ; greenflow, ; stojkovic2024dynamollm, ; samsi2023words, ; nguyen2024towards, ). We anticipate future work on energy efficiency of co-locating side tasks with LLM training.
@@ 9. Conclusion
We propose FreeRide, a middleware system that bridges the gap between the available yet hard-to-utilize bubbles in pipeline parallelism and running generic GPU side tasks to harvest them. It provides programming interfaces that abstract the life cycle of a side task as different states of a state machine and allows programmers to implement side tasks with little engineering effort. The side task manager and side task workers manage bubbles and side tasks, and reduce the performance overhead of side tasks on pipeline training. Our evaluation shows that, on average, FreeRide achieves 7.8% cost savings for long-running and expensive pipeline training with a negligible performance overhead of about 1%.
Acknowledgements.
We thank the anonymous reviewers and our shepherd Ruben Mayer for their constructive feedback to improve this paper. This work was supported by the Natural Sciences and Engineering Research Council of Canada (NSERC) and the Undergraduate Research Assistantship program of the Cheriton School of Computer Science at the University of Waterloo.